OpenStudy (loser66):

if they say: If the block starts from rest 12.0 m up the plane from its base, what will be ..... where is the 12.0m leg?

OpenStudy (loser66):



OpenStudy (loser66):

@e.mccormick

OpenStudy (loser66):

AB or AC?

OpenStudy (loser66):

Problem 48

OpenStudy (e.mccormick):

I think it is BC, but I have not found a single reference.

OpenStudy (loser66):

you mean you are not sure?

OpenStudy (e.mccormick):

Never took physics.

OpenStudy (loser66):

that's why I don't ask you physics, I ask you English. right? hehehe.

OpenStudy (e.mccormick):

Yep. Or some of computer science. Hehe.

OpenStudy (loser66):

wait for me, I will not let you free next semester when I take computer science.

OpenStudy (e.mccormick):

Hehe. I'll be sure to forget it all by then.

OpenStudy (ybarrap):

"The block starts from rest 120 m up the plane from it's base" means that the block is perpendicular to the floor (i.e. the ground plane, where it's base is located). So \(\overline{AB}=120~m\).

OpenStudy (loser66):

@ybarrap so, I have to use tan \(\theta\) to find out AC and then apply AC = 1/2at^2 as it moves horizontal direction, ignore BC, right?

OpenStudy (ybarrap):

I meant 12.0 m, not 120 m. AB = 1/2 at^2. The vertical drop will be like free fall with nothing to slow it down, like friction.

OpenStudy (e.mccormick):

Never would have guessed AB since they define "the plane" as BC. (II) The block shown in Fig. 4-43 has mass m = 7.0 kg and lies on a fixed smooth frictionless plane tilted at an angle \(\theta = 22.0^\circ\) to the horizontal. (a) Determine the acceleration of the block as it slides down the plane. (b) If the block starts from rest 12.0 m up the plane from its base, what will be the block's speed when it reaches the bottom of the incline?

OpenStudy (ybarrap):

I think you are right, I did not read that part carefully. It slides "down the plane" and it is 12.0 m "up the plane" @e.mccormick makes sense. I would agree.

OpenStudy (ybarrap):

I still think AB = 1/2 at^2. The vertical drop will be like free fall with nothing to slow it down, like friction.

OpenStudy (loser66):

@ybarrap I understand what you mean. the time for the object go from B to C is calculated by vertical direction which is the same if we calculate via horizontal direction. I just want to figure out what is 12.0 m length. Obviously, if I go backward and say AC =12, my answer goes wrong.

OpenStudy (ybarrap):

No, we think that BC=12.0 m.

OpenStudy (loser66):

whaaatt? you confused me. why BC?

OpenStudy (loser66):

oh, oh, I got you. heheh. sorry friend

OpenStudy (ybarrap):

No problem. The question is not very clear, but that BC = 12.0 makes the most sense.

OpenStudy (loser66):

got you, thank you